OpenStudy (anonymous):

Which of the points is a solution to -5x-3y=15? my answers are: (-5,0) (-3,0) (1,3) (3,0) plz someone help me!!!

OpenStudy (the_fizicx99):

B

OpenStudy (anonymous):

your answer should be (-3,0) and (0,-5)

OpenStudy (anonymous):

so its -3,0

OpenStudy (anonymous):

yeah..:)

OpenStudy (the_fizicx99):

-5x -3y = 15 (x,y) --> (-3,0) -5(3) -3(0) = 15 15 + 0 = 15 15 = 15 so yesh